OpenStudy (anonymous):

Condense the expression: log (base 3) 2x + 3 log (base 3) 4x Please show work! Thank you ^_^

OpenStudy (anonymous):

\[\log_{3} (4x(2x+3))\] \[\log_{a} MN=\log_{a}M+\log_{a}N\]

OpenStudy (anonymous):

\[\log_{2} (x+8)=4\]Thanks :) how do i do this one?

OpenStudy (anonymous):

whoa luis don't know what you did there but that's not right

OpenStudy (anonymous):

it's right; that is the answer my teacher got.

OpenStudy (anonymous):

all the work is right except how u got to the first line

OpenStudy (anonymous):

oh i didn't see the 3 before the log in the 2nd part, my bad, you are right

OpenStudy (anonymous):

log base 2 ? On both sides?

OpenStudy (anonymous):

for the 2nd one, put it into exponential form. 2^4=x+8

OpenStudy (anonymous):

Thank you. (: